OpenStudy (anonymous):

Write a recursive formula for 16, 9, 7, 2, 5....

OpenStudy (anonymous):

what is 16-9?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

What's the term after 9 in the sequence?

OpenStudy (anonymous):

What is your hypothesis?

OpenStudy (anonymous):

9-2=7 an 7-2=5

OpenStudy (anonymous):

Hypothesis?

OpenStudy (anonymous):

ok but how write the formula?

OpenStudy (turingtest):

name the terms in your head

OpenStudy (turingtest):

what is the relationship between successive terms?

OpenStudy (anonymous):

when you subtruct the first number by the second you will get the third and so on

OpenStudy (turingtest):

exactly the first two terms are just given after that the following terms are generated according to the pattern you just identified so call some number past the third number\[a_n\]how can we relate that number to the numbers preceding it\[a_{n-1},a_{n-2}\]?

OpenStudy (anonymous):

an-7,an-2,an-5

OpenStudy (turingtest):

I think you are misunderstanding the notation...

OpenStudy (turingtest):

\[a_{n}\]is the nth number in the sequence\[16, 9, 7, 2, 5... \]\[a_1,a_2,a_3,a_4,a_5...\]so the number in the subscript is just telling you what term we are on now, what is the relationship between\[a_5,a_4,\text{ and }a_3\]?

OpenStudy (anonymous):

subtracting 1

OpenStudy (turingtest):

write out a formula for \(a_5\) in terms of \(a_4\) and \(a_3\)

OpenStudy (anonymous):

how do I write a formula

OpenStudy (turingtest):

try it with number let me make sure you understand what I am asking... what is \(a_5\) ?

OpenStudy (anonymous):

5

OpenStudy (turingtest):

yes, what is \(a_4\) ? what is \(a_3\) ? how are these three numbers ralated?

OpenStudy (turingtest):

related*

OpenStudy (anonymous):

4,3,2 they are decreasing as you go along by 1

OpenStudy (turingtest):

no you have misunderstood\[a_1=16\]\[a_2=9\]\[a_3=7\]\[a_4=2\]\[a_5=5\]\[...\]the numbers in the bottom right (called subscripts) just tell you the term we are on

OpenStudy (anonymous):

and will that be considered a formula

OpenStudy (turingtest):

so in general when we write \[a_n=n^{th}\text{ term }\]to write a formula we need to generalize the relationship between any three numbers in the list, so no, listing out the given terms is how we start, not the answer

OpenStudy (turingtest):

what is \(a_3\) ?

OpenStudy (anonymous):

3

OpenStudy (turingtest):

is it? what is the third term on the list?

OpenStudy (anonymous):

7

OpenStudy (turingtest):

yes what is \(a_2\) ?

OpenStudy (anonymous):

9

OpenStudy (turingtest):

yes what is \(a_1\) ?

OpenStudy (anonymous):

16

OpenStudy (turingtest):

now you're getting the idea now what is the relationship between those three numbers?

OpenStudy (anonymous):

subtracting the first 2 to get the 3rd number

OpenStudy (turingtest):

right now write that out mathematically (if you still can't do it with an then just write it out with the numbers....)

OpenStudy (anonymous):

16-9=7,7-2=5

OpenStudy (turingtest):

excellent now all you need to do is rewrite this with the \(a_n\) notation and the formula comes out instead of writing 16-9=7 write a_?-a_?=a_? fill the ?'s with the correct subscript (number of the term)

OpenStudy (anonymous):

a-16=9=a-9=7=a-7=2=a-5

OpenStudy (turingtest):

no, remember what \(a_1\) means what is \(a_1\) ?

OpenStudy (turingtest):

if you can't write this using the equation editor write it as a_1=

OpenStudy (anonymous):

16-a=9,9-a=7,7-a=2,a-2=5

OpenStudy (turingtest):

what is a_1 ?

OpenStudy (anonymous):

16

OpenStudy (turingtest):

what is a_2 ?

OpenStudy (anonymous):

9

OpenStudy (turingtest):

what is a_3 ?

OpenStudy (anonymous):

7

OpenStudy (turingtest):

now you showed that 16-9=7 so rewrite the above expression with the a_n notation a_?-a_?=a_? (there should only be three terms)

OpenStudy (anonymous):

a16-a9=a7

OpenStudy (turingtest):

16 is not a16 16 is a_1

OpenStudy (turingtest):

a16 would be the 16th term of the sequence (we don't know what that is yet) you are over complicating the situation. all I want you to do is recognize the terms which you just did a moment ago

OpenStudy (turingtest):

fill in the other two numbers 9=a_? (it's \(not\) a_9, it's the number of the \(term\))

OpenStudy (anonymous):

a_1-a_2=7 a_3_a_4=2a_3-a_4=5

OpenStudy (turingtest):

almost, but please don't write it all in a line like that, it is hard to read.... if you started writing a_1-a_2=7 that is good, but you still need to sub for 7=a_?

OpenStudy (anonymous):

what do you mean couldn't write like the way I wrote

OpenStudy (turingtest):

did you try to write a_1-a_2=7 a_3-a_4=2 a_3-a_4=5 ??? if so that is not quite right, we want no numbers left just call everything by the term name

OpenStudy (turingtest):

so what is term is 7 (a_(what number??)) a_1-a_2=a_?

OpenStudy (anonymous):

3 term

OpenStudy (turingtest):

so we can write a_1-a_2=a_3 understand??

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

ok for the sake of simplicity let's turn this around a_3=a_1-a_2 now what about the other terms? a_4=? a_5=?

OpenStudy (anonymous):

a_4=a_2-a_3

OpenStudy (turingtest):

nice!

OpenStudy (anonymous):

a_5=a_3-a_4

OpenStudy (turingtest):

ok, so now you need to find a way to write this in general for a_n=?

OpenStudy (anonymous):

whats n

OpenStudy (turingtest):

to do this try calling n=5 what is the formula then? \[a_5=a_3-a_4\]let \(n=5\) \[a_n=a_?-a_?\]that is, how are the three subscripts related (if n=5 then 3=n-?)

OpenStudy (turingtest):

n is the unknown n'th term in the sequence

OpenStudy (turingtest):

that is, whatever term n they want will specified by our formula

OpenStudy (anonymous):

ok so how do i write the formula?

OpenStudy (turingtest):

like I say, try it with any three numbers\[a_4=a_2-a_3\]if we rename \(4=n\) how would we write 3 in terms of n ? (as 3=n-(something) ) i.e. 3=4-x